OpenStudy (lonnie2018):

Put 3x – 2y = 12 in slope-intercept equation?

OpenStudy (blacksteel):

Slope-intercept equations take the form of y = mx + b, where m is the slope and b is the y-intercept. So let's take our starting equation: 3x - 2y = 12 The first step is to get isolate the y variable: -2y = 12 - 3x Now we want to divide both sides by the coefficient of y: -2y/-2 = (12 - 3x)/-2 y = (3/2)x - 6 This is of the form y = mx + b, so we're done. The slope is 3/2, and the y-intercept is -6.